# PEVLM: Parallel Encoding for Vision-Language Models

### v1

@@ 4.2 Performance Evaluation
In this section, we evaluate the performance of PEVLM. To facilitate experimentation, we built PEVLM on the SGLang framework, which is a fast serving framework for large language models and vision-language models, widely used in cloud production deployment. And to better reflect practical deployment scenarios, we designed the experiments from two perspectives. (i) We first test with workloads of approximately 100k tokens to assess the acceleration performance of PEVLM in cloud computing environments. (ii) Furthermore, we evaluate PEVLM on the LongVideoBench dataset under a given latency constraint, in order to measure its effectiveness in edge scenarios where both computational resources and latency are limited.
4.2.1 Setup
@@
Since the LLM component in VLM models remains largely consistent across different implementations, performance differences arise primarily from hyperparameter configurations rather than model weights. Consequently, we conducted computational efficiency evaluations on the Qwen2.5-VL model. The primary objective of this experiment is to assess the improvements in computational efficiency introduced by PEVLM in computational efficiency. All experiments were carried out on an H20-96G machine. Given that PEVLM primarily optimizes the attention mechanism within the LLM, we measured the execution time of both the entire LLM module and its individual attention layers separately.
Figure 4: We evaluated the performance of PEVLM under different configurations of (sink_size / block_size). Using the Qwen2.5-VL-7B-Instruct model, we tested its LLM and Attention time on a video dataset with an average duration of 10 minutes (approximately 106k tokens per sample). "Attn Time" designates the Attention Computation time, whereas "Other Time" signifies the aggregate of all remaining computational costs within the LLM excluding Attention Computation. The "Attn SpeedUP" represents the performance acceleration of PEVLM relative to the "Full Attn". The "LLM SpeedUP" denotes the overall performance acceleration of the LLM in PEVLM compared to the "Full Attn".
4.2.2 Results
@@
As illustrated in Figure 4, the end-to-end inference speed of the LLM is improved by 2.58× to 3.14×. Notably, the attention module achieves a speedup ranging from 7.79× to 33.81×, while the execution time of non-attention components remains largely unaffected. As the block size or sink size increases, the acceleration effect of the attention module decreases significantly. This observation is consistent with the analysis presented in the Appendix. Moreover, because non-attention components constitute a substantial portion of the total runtime in PEVLM, the overall end-to-end speedup of the LLM moderately decreases from 3.14× to 2.58×.
4.3 Accuracy with Limited Latency
@@
4.3.1 Setup
@@
We use LongVideoBench as the evaluation benchmark and introduce a latency constraint during inference: any sample whose inference time exceeds the predefined latency threshold is considered a failure. The experiments are conducted on the H20-96B platform. We select Qwen2.5-VL-7B-Instruct with Full-Attention as the baseline, and compare several methods, including Full-Attention, Block Attention, APE, Star Attention, PEVLM, and using smaller models (Qwen2.5-VL-3B-Instruct). Since the other methods are finetune-free, we did not finetune the weights in the Block Attention test. Although position encoding does not affect inference lantency, reusing position encoding performs extremely poorly accuracy on the Qwen2.5-VL model. Therefore, in the APE and Star Attention tests, we adopted the method of retaining Full-Attention’s position encoding.
4.3.2 Results
Table 2: LongVideoBench Accuracy with Limited Latency
method sink(anchor) block No Limit $\uparrow$ 40s $\uparrow$ 30s $\uparrow$ 20s $\uparrow$
Full Attn - - 60.43% 59.69% 27.08% 23.26%
Full Attn (3B) - - 55.05% 55.05% 55.05% 24.38%
Block Attn 0 4096 54.00% 54.00% 53.78% 53.55%
APE 256 4096 59.91% 59.91% 59.91% 59.61%
Star Attn 4096 4096 61.41% 61.41% 61.41% 28.20%
PEVLM sys+2frames 16frames 61.18% 61.18% 61.18% 61.03%
PEVLM sys+16frames 16frames 62.23% 62.23% 62.23% 60.28%
@@ 4.3.2 Results
As shown in Table 2, the accuracy of Full-Attention drops significantly when the latency constraint is reduced from 40s to 30s, while the accuracy of Full-Attention 3B also declines sharply when the latency constraint is tightened from 30s to 20s. Due to its larger sink and block sizes, StarAttn gets more OPs and exhibits a noticeable accuracy drop between 30s and 20s. In contrast, Block Attention, APE, and PEVLM maintain relatively high accuracy even at 20s. Notably, although larger sink and block sizes (e.g., sys+16frames, block=16frames) achieve higher accuracy under looser latency constraints, the smaller sink size (sink=sys+2frames, block=16frames) delivers better accuracy at the 20s latency limit due to its lower latency.
5 How does each component in PEVLM contribute to the performance?
5.1 Split The Video by Frames instead of Tokens
Due to the inherent nature of videos being composed of multiple frames, where intra-frame information captures spatial details at a single time point, and inter-frame information represents temporal dynamics, ignoring this structure and arbitrarily dividing blocks based on token count may compromise the spatial integrity at critical temporal boundaries. Therefore, we choose to divide the video into blocks based on frames. This design is supported by the comparison between Star Attention with Full-Attention positional encoding and PEVLM, as shown in Table 1. Across different models, frame-based block partitioning consistently provides accuracy gains, albeit varying in magnitude, over token-based partitioning.
5.2 Sink Size
@@
As analyzed in Section 3.2.2, the inference latency of PEVLM increases quadratically with the sink size, which is consistent with the experimental results shown in Figure 4.
@@
Furthermore, we investigate the impact of sink size on model accuracy. As observed in Section 2.1, it is essential to involve the initial frames of the video in the sink. This observation is validated in Figure 5, where incorporating the initial frames into the sink leads to varying degrees of accuracy improvement across different models.
@@
Moreover, as shown in Figure 5, the accuracy of Qwen2.5-VL and LLaVA-Video shows an upward trend as the sink size increases, though the improvement remains marginal. In contrast, LongVILA exhibits fluctuations in accuracy with varying sink sizes but does not show a consistent upward or downward trend. However, because of the quadratic increase in latency, using longer sink sizes is not favorable for performance. Therefore, it may be necessary to conduct a simple test prior to deployment to determine an appropriate sink size. For the three models shown in Figure 5, setting the sink size between 14 and 18 already yields strong accuracy. If latency constraints are also taken into account, this could lead to a trade-off between performance and accuracy. For example, as demonstrated in the results of Section 4.3, under a 20-second latency constraint, using only sys+2frames as the Sink Block achieves better accuracy than using sys+16frames.
Figure 5: The impact of sink/block sizes on accuracy
5.3 Block Size
@@
As analyzed in Section 3.2.2, the inference latency of PEVLM also increases quadratically with the block size as well as the sink size, which is consistent with the experimental results shown in Figure 4. However, the block size has a more significant impact on performance because the coefficient of block-size’s quadratic term in Equation 10 is much larger than that of the sink size.
@@
5.4 Position Encoding
As analyzed in Section 2.1, using Full-Attention position embeddings, rather than reusing position embeddings across blocks, should yield attention distributions that more closely resemble those of Full-Attention, thereby aligning better with the model’s intended behavior. This is validated by the experiments presented in Section 4.1. As shown in Table 1, for several parallel-encoding-based methods (Block-Attention, APE, Star Attention, and PEVLM), using Full-Attention position embeddings significantly improves accuracy in both Qwen2.5-VL and LongVILA. In contrast, for the LLaVA-Video model, the benefit is marginal, likely because its attention weight distribution remains relatively stable even when using parallel encoding.
6 Conclusion
This study investigates the potential of parallel-encoding-based methods in VLMs, which can reduce the computational complexity of attention mechanisms for fast inference and reuse positions for long-context processing but often result in performance degradation. To address this issue, we propose PEVLM, a training-free method designed to enable accurate, fast, and long-context video processing. PEVLM achieves this by aligning the attention weight distribution of parallel encoding with Full-Attention through two key steps: selecting an appropriate sink size and block size, and utilizing Full-Attention position embeddings. Empirically, we demonstrate that PEVLM enhances both accuracy and efficiency in long video processing tasks while successfully scaling to process hundreds of chunks in parallel across various settings.
7 Future Work
@@
In autonomous driving scenarios, a common strategy is to use several historical frames as input to make decisions for the next action. Due to the computational nature of the attention mechanism, most current approaches require re-computation of all historical frames as the temporal step advances. Building upon the current PEVLM framework, we propose integrating the rotating positional embedding method from Block-Attention (ma2025blockattentionefficientprefilling, ), which enables the reuse of the KVCache from all historical frames preceding the current timestep. This approach has the potential to yield an order-of-magnitude improvement in efficiency for such applications.
Many of the existing methods based on parallel-encoding, including PEVLM, can achieve accuracy that exceeds Full-Attention. This phenomenon is attributed to the fact that parallel-encoding-based methods can mitigate the "lost in the middle" problem, thereby enabling them to achieve accuracy surpassing Full-Attention when processing long contexts (yang2025ape, ). The reason why parallel-encoding can mitigate the "lost in the middle" issue is hypothesized to be due to its shortening of the distance between the context and the query. However, more concrete evidence is needed to support this claim.

### v2

@@ Introduction
In this section, we test PEVLM from multiple perspectives against several different VLMs. The primary objective is to evaluate the accuracy, computational efficiency, and real-world applicability of PEVLM for long-video processing. And to better reflect practical deployment scenarios, we designed the experiments from two perspectives. (i) We first test with workloads of approximately 100k tokens to assess the acceleration performance of PEVLM in cloud serving platforms. (ii) Furthermore, we evaluate PEVLM on the LongVideoBench dataset under a given latency constraint, in order to measure its effectiveness in edge scenarios where both computational resources and latency are limited.
Accuracy Evaluation
Setup
The main objective of this experiment is to analyze the impact of PEVLM on accuracy in long-video understanding tasks. We have chosen the LongVideoBench (Zhang and Wang 2024), VideoMME (Fu et al. 2024), EgoSchema (Mangalam, Akshulakov, and Malik 2023), and MVBench (Li et al. 2024a) as benchmarks, which are designed to comprehensively evaluate multimodal models in video understanding. We have carried out experiments using LLaVA-Video (Zhang et al. 2024), LongVILA (Chen et al. 2024), and Qwen2.5-VL (Bai et al. 2025) as test models. To minimize any adverse effects on accuracy, all evaluations were performed using bf16 precision.
As shown in Table 1, we compared the accuracy of Full-Attention, Block-Attention, APE, Star-Attention, and PEVLM in the datasets. Since the other methods are finetune-free, we did not finetune the weights in the Block Attention test. For APE, to accommodate the needs of large-scale deployment, we tested with T=0. Regarding the sink block (shared prefix), we selected the system prompt by referring to the code provided by APE. For Star-Attention, we implemented its equivalent algorithm on a single node, with the anchor size set equal to the block size, as the configuration in the Block-Attention paper (Ma, Wang, and Tian 2025). The block size was uniformly set to 4096 across all algorithms. Smaller block sizes resulted in reduced accuracy for all methods, while larger blocks significantly degraded inference performance. For PEVLM, we set the sink size and block size to about 16 frames, which corresponds to approximately 4k tokens for Qwen2.5-VL and LongVILA, and about 3k tokens for LLaVA-Video.
Results
The experiments show that PEVLM achieves an accuracy comparable to, and in some cases even surpassing, that of Full-Attention across all three evaluated models. Both APE and Star-Attention also approach or exceed the accuracy of Full-Attention when Full-Attention’s position encoding is retained. However, Block-Attention exhibits a notable accuracy drop compared to Full-Attention in most settings, primarily due to the lack of alignment-specific fine-tuning.
Although parallel encoding is generally expected to degrade performance, it can instead improve accuracy in long-context scenarios. This is because parallel encoding effectively shortens the total token length involved in each softmax computation, thereby mitigating the softmax feature degradation that occurs with long contexts (Veličković et al. 2025). As a result, the softmax output more robustly approximates an ideal sharp function, which ultimately enhances the performance of some models under long-context conditions.
Performance Evaluation
In this section, we evaluate the performance of PEVLM. To facilitate experimentation, we built PEVLM basied on the SGLang, which is a fast serving framework for large language models and vision-language models, widely used in cloud production deployment.
Figure 4: We evaluate PEVLM under different block_size settings. The sink_size is set to (system prompt+one frame). "Attn Time" refers to attention computation time, and "Other Time" covers all other LLM costs. "Attn SpeedUP" and "LLM SpeedUP" indicate the acceleration over Attention layer and overall LLM performance.
Setup
@@ Results
Since the LLM component in VLM models remains largely consistent across different implementations, performance differences arise primarily from hyperparameter configurations rather than model weights. Consequently, we conducted computational efficiency evaluations on the Qwen2.5-VL model. The primary objective of this experiment is to assess the improvements in computational efficiency introduced by PEVLM. All experiments were carried out on an NVIDIA H20-96G GPGPU. Given that PEVLM primarily optimizes the attention mechanism within the LLM, we measured the execution time of both the entire LLM module and its individual attention layers separately.
Figure 5: We evaluate PEVLM under different sink_size settings. The block_size is set to (16 frames). "Attn Time" refers to attention computation time, and "Other Time" covers all other LLM costs. "Attn SpeedUP" and "LLM SpeedUP" indicate the acceleration over Attention layer and overall LLM performance.
Results
@@
As illustrated in Figure 4 and Figure 5, the end-to-end inference speed of the LLM is improved by 2.58× to 3.14×. Notably, the attention module achieves a speedup ranging from 7.79× to 33.81×, while the execution time of non-attention components remains largely unaffected. As the block size or sink size increases, the acceleration effect of the attention module decreases significantly. This observation is consistent with the Equation 11. Moreover, because non-attention components constitute a substantial portion of the total runtime in PEVLM, the overall end-to-end speedup of the LLM moderately decreases from 3.14× to 2.58×.
Accuracy with Limited Latency
@@
Table 2: Accuracy with Limited Latency
Full Attn Full Attn PEVLM
sink - - sys+2f sys+16f
block - - 16f 16f
No Limit 60.43% 55.05% 61.18% 62.23%
40s 59.69% 55.05% 61.18% 62.23%
30s 27.08% 55.05% 61.18% 62.23%
20s 23.26% 24.38% 61.03% 60.28%
Setup
@@
We use LongVideoBench as the evaluation benchmark and introduce a latency constraint during inference: any sample whose inference time exceeds the predefined latency threshold is considered a failure. The experiments are conducted on the H20-96B platform. We select Qwen2.5-VL-7B-Instruct with Full-Attention as the baseline, and compare Full-Attention, PEVLM, and using smaller models (Qwen2.5-VL-3B-Instruct).
Results
@@
As shown in Table 2, the accuracy of Full-Attention drops significantly when the latency constraint is reduced from 40s to 30s, while the accuracy of Full-Attention 3B also declines sharply when the latency constraint is tightened from 30s to 20s. PEVLM maintain relatively high accuracy even at 20s. Notably, although larger sink and block sizes (e.g., sys+16frames, block=16frames) achieve higher accuracy under looser latency constraints, the smaller sink size (sink=sys+2frames, block=16frames) delivers better accuracy at the 20s latency limit due to its lower latency.
How does each component in PEVLM contribute to the performance?
In Table 3, we present an ablation study to evaluate the contribution of each component in PEVLM, including the use of Sequential Position Embedding (SP), Dividing the Video by Frames (DF), and Adding Frames to the Sink (FS). The results are averaged across the four base models reported in Table 1. We observe that sequential parallel encoding yields the most significant improvement in accuracy, achieving an average performance gain of 12.64%. Incorporating the initial frames into the sink block leads to an additional average accuracy gain of 1.54%. In contrast, dividing the video by frames provides only a marginal improvement compared to directly segmenting the video without this strategy.
Table 3: Ablation study of PEVLM components on LongVideoBench
SP DF FS LongVideoBench VideoMME
46.67% 45.31%
$\checkmark$ 56.75% 60.51%
$\checkmark$ $\checkmark$ 56.77% 60.63%
$\checkmark$ $\checkmark$ $\checkmark$ 58.26% 62.22%
Position Encoding
We further analyze the component with the greatest impact on accuracy: the use of sequential position embeddings. As discussed before, adopting sequential position embeddings instead of reusing position embeddings across blocks produces attention distributions that more closely resemble those of Full-Attention, thereby aligning better with the model’s expected behavior. As shown in Table 4, the use of sequential position embeddings significantly improves accuracy in both Qwen2.5-VL and LongVILA. In contrast, for the LLaVA-Video model, the benefit is marginal, likely because its attention weight distribution (as shown in Figure 9 in the Appendix) remains relatively stable even when using parallel encoding.
Table 4: Accuracy $\uparrow$ on LongVideoBench dataset with Different Position Embedding Strategy
model method reuse pos sequential
Qwen2.5-VL- APE (T=1.0) 30.44% 59.39%
7B-Instruct Star Attn 29.54% 61.41%
PEVLM 14.73% 62.23%
APE (T=1.0) 51.46% 53.40%
7B-256f Star Attn 51.76% 53.63%
PEVLM 49.96% 53.70%
LLaVA-Video- APE (T=1.0) 58.19% 58.12%
7B-Qwen2 Star Attn 58.71% 58.49%
(128) PEVLM 58.34% 59.01%
Sink & Block Size
@@
As analyzed in the third section, the inference latency of PEVLM increases quadratically with the sink size, which is consistent with the experimental results shown in Figure 5.
@@
Furthermore, we investigate the impact of sink size on model accuracy. As observed in the second section, it is essential to involve the initial frames of the video in the sink. This observation is validated in Figure 6, where incorporating the initial frames into the sink leads to varying degrees of accuracy improvement across different models.
Figure 6: The impact of sink sizes on accuracy
@@
Moreover, as shown in Figure 6, the accuracy of all models shows an upward trend as the sink size increases. However, because of the quadratic increase in latency, using longer sink sizes is not favorable for performance. Therefore, it may be necessary to conduct a simple test prior to deployment to determine an appropriate sink size. For the three models shown in Figure 6, setting the sink size between 14 and 18 already yields strong accuracy. If latency constraints are also taken into account, this could lead to a trade-off between performance and accuracy. For example, as demonstrated in the results of Table 2, under a 20-second latency constraint, using only sys+2frames as the Sink Block achieves better accuracy than using sys+16frames.
Figure 7: The impact of block sizes on accuracy
@@
As shown in Figure 4, the inference latency of PEVLM also increases quadratically with block size, which is consistent with the experimental results shown in Figure 4. However, the block size has a more significant impact on performance because the coefficient of the block size’s quadratic term in Equation 11 is much larger than that of the sink size.
@@
Conclusion
This study investigates the potential of parallel-encoding-based methods in VLMs, which can reduce the computational complexity of attention mechanisms for fast inference and reuse positions for long-context processing but often result in performance degradation. To address this issue, we propose PEVLM, a training-free method designed to enable accurate, fast, and long-context video processing. PEVLM achieves this by aligning the attention weight distribution of parallel encoding with Full-Attention through two key steps: selecting an appropriate sink size and block size, and utilizing sequential position embeddings. Empirically, we demonstrate that PEVLM enhances both accuracy and efficiency in long video processing tasks while successfully scaling to process hundreds of chunks in parallel across various settings.
Future Work
@@ Conclusion
Beyond video inputs, an important future direction is extending PEVLM to handle richer multimodal inputs such as LiDAR, map priors, and inertial measurements, which are commonly used in domains like autonomous driving and robotics. These modalities often exhibit different spatial and temporal characteristics compared to visual tokens, posing unique alignment and fusion challenges for attention-based models. Enhancing PEVLM with cross-modal position encoding strategies and adaptive sink block selection could further improve its generalization to complex, heterogeneous input streams.
One promising direction is extending PEVLM to streaming or online scenarios, such as autonomous driving, where the model continuously receives multimodal inputs (e.g., video frames and LiDAR) over time. In these settings, a fixed-length sliding temporal window is often adopted: new data is appended while the oldest is discarded, maintaining a constant context size that reflects the most recent few seconds or minutes of input. If PEVLM can be adapted to such settings, it would allow the model to compute only over the newly arrived data at each inference step, rather than reprocessing the entire history. This could lead to significant performance gains—potentially improving inference speed by several times or even orders of magnitude. Such an extension would enable real-time video-language reasoning under strict latency constraints and dynamic, continuously evolving environments.
Appendix A Benchmark Details
We evaluate our method on several video understanding benchmarks that test different aspects of video comprehension:
EgoSchema
EgoSchema (Mangalam, Akshulakov, and Malik 2023) is a large-scale benchmark designed to evaluate Multimodal Large Language Models (MLLMs) on egocentric video understanding. It consists of 100 hours of first-person video data spanning 1,270 daily activity episodes across diverse real-world environments. The benchmark introduces over 10,000 manually curated question-answer pairs, covering tasks such as object grounding, human-object interaction, activity reasoning, and intent prediction.
MVBench
MVBench (Li et al. 2024a) is a comprehensive benchmark designed to evaluate Multimodal Large Language Models (MLLMs) on multi-granular video understanding. It consists of 5 task categories—including moment-level, frame-level, clip-level, video-level, and holistic video understanding—covering a wide range of temporal scopes and reasoning demands. The benchmark includes 2,562 manually annotated questions grounded in 4,119 diverse video clips, selected from real-world scenarios.
Each question is carefully designed to probe different levels of spatiotemporal understanding, from fine-grained object recognition and short-term motion tracking to long-term event inference. MVBench offers a unified and challenging evaluation protocol to assess the generalization and reasoning ability of MLLMs across granularities.
Unlike conventional third-person video datasets, EgoSchema emphasizes embodied perception and temporal reasoning from an egocentric perspective, posing unique challenges for spatial understanding, long-term memory, and causal inference in MLLMs.
Video-MME
Video-MME (Fu et al. 2024) is a comprehensive evaluation benchmark for assessing the video understanding capabilities of Multimodal Large Language Models (MLLMs). It spans 6 primary visual domains and 30 subfields, covering a diverse range of video types and temporal scenarios. The benchmark includes 900 videos with durations ranging from 11 seconds to 1 hour, totaling 254 hours of content.
To evaluate fine-grained visual and temporal reasoning, 2,700 manually annotated question-answer pairs are provided. Video-MME challenges models to comprehend both short and long video clips across different temporal granularities, making it a rigorous benchmark for evaluating the core video processing capabilities of MLLMs.
LongVideoBench
LongVideoBench (Zhang and Wang 2024) is a benchmark specifically designed to evaluate the long-context video understanding capabilities of Multimodal Large Language Models (MLLMs). It features 1,760 videos spanning 12 diverse real-world scenarios, with video durations ranging from 5 minutes to 2 hours, totaling over 1,000 hours of content. The benchmark includes 2,400 manually annotated multi-choice questions targeting key aspects of long video comprehension.
LongVideoBench focuses on challenging long-range temporal reasoning, event tracking, and global understanding across extended video content. It aims to assess whether models can maintain coherence, memory, and attention over prolonged contexts, making it a rigorous testbed for long-form video modeling.
Appendix B Attention Weights Distribution of LLaVA-Video and LongVILA
(a) Full-Attention
(b) Reuse position embeddings
(c) Sequential position embeddings
Figure 8: Attention Weight Distributions of LongVILA
(a) Full-Attention
(b) Reuse position embeddings
(c) Sequential position embeddings
Figure 9: Attention Weight Distributions of LLaVA-Video